OpenStudy (stamp):

Simplify the following expression by using the double-angle formula. (Expression inside)

OpenStudy (stamp):

\[cos^25ø-sin^25ø\]

OpenStudy (stamp):

cos(10x) ???

OpenStudy (anonymous):

Recall that \(\large \cos^2\alpha - \sin^2\alpha = \cos(2\alpha)\). In this case, \(\large\alpha = 5\theta\), so what do you get as the simplified trig function?

OpenStudy (stamp):

Well, cos(10ø) to be consistent?

OpenStudy (anonymous):

And I see you answered it as I was writing up my answer. You are correct. :-)

OpenStudy (stamp):

Thank you.